# Scalable Qualitative Coding with LLMs: Chain-of-Thought Reasoning Matches Human Performance in Some Hermeneutic Tasks thanks: Code and data available at https://osf.io/k4fg9

### 26 Jan 2024 v1

@@ 1.1 Automating Content Analysis: Past and Present
Early studies of content analysis with LLMs are encouraging. Xiao et al.[9] demonstrate moderate success, Cohen’s $\kappa=0.61$ and $\kappa=0.38$, in two linguistic tasks using GPT-3. Chew et al.[10] report high success on many of 19 tasks across three datasets, and results that are indistinguishable from random for others. It is difficult to evaluate their results due to the choice of Gwet’s AC1, which is biased toward agreement on negative codings rather than positive, whereas most standard measures of intercoder reliability do the opposite[24]. However, we commend Chew et al.’s approach to adapting codebooks for LLMs, which is communicated with great detail and clarity.
We present here three advances to these studies. 1) We report the first results of automating qualitative coding using GPT-4, which, along with other recent models, greatly improves upon many of GPT-3’s capabilities[25, 26, 19]. 2) We provide the first conclusive evidence that LLMs are capable of human-equivalent performance in qualitative coding, and do so on larger passages of text, wherein meaning is often woven through multiple interrelated clauses. 3) We demonstrate that GPT is better at interpreting text when it is tasked with justifying its coding decisions (chain-of-thought prompting) rather than applying codes without an accompanying explanation.
@@ 2.1 Adapting a Codebook for an LLM
Excerpt 1 demonstrates an effective method of prompting GPT to provide rationale for its code selections.
Excerpt 1: Prompting GPT for rationale for code application.
When you evaluate the passage, provide a justification of why you did or did not apply the code.
Then list the code in the following fashion if you applied the code:
Justification: [insert 2-3 sentence reasoning for applying the code here]
Codes Applied:
      - Scholar
Otherwise you can format it like this:
Justification: [insert 2-3 sentence reasoning for not applying the code here]
Codes Applied:
 - None
Do not write anything in your reply after listing the ‘‘Codes Applied:’’
@@
Human coders do not rely solely on a written codebook. Their interpretation of the codes is enriched through the codebook development process, discussions with fellow coders, and supplementary oral instructions. An LLM lacks this interactive and historical context and must interpret codes entirely from written descriptions. Our work modifying the codebook for GPT revealed information that, while implicitly understood by the code developers, wasn’t explicitly stated in the code descriptions. This process not only aided in refining the codebook for automated coding, but also improved our own understandings of the codes. This ultimately led to clearer definition of the codes, thereby enhancing future manual coding processes as well.
Often, we encountered cases where ambiguous phrasing was obvious to humans, but challenging for the LLM. Our codebook contains two codes that relate to Du Bois’s reputation among academics and activists. These codes are meant to evaluate whether Du Bois appears in a news story because either an academic or activist mentioned him. Initially, we titled this code “Academic Repute,” which worked well for human coders. GPT, however, consistently misinterpretted this code as pertaining to Du Bois’s esteem as an academic, rather than among or by academics (the meaning of “among” remains ambiguous even here). We tried numerous iterations this code without success. Nevertheless, altering the title of the code to the far more literal “Out of the Mouth of Academics” dramatically improved performance, even when paired with the original code description. In another case, the code titled “Social/Political Activism” was revised to “Social/Political Advocacy” because GPT did not consider social critique to be a form of activism, even when it was specifically instructed to.
We found that words indicating how much the model should draw on context or its own outside knowledge had large impacts on the model’s outputs, often to the desired effect. In particular, instructing the model to restrict itself to “explicit” meanings, or to draw on “implicit” meanings, often helped the model with part of a code description it had struggled with. Both mandatory (do) and prohibitory (do not) phrasing were observed by the model, though mandatory phrasing seemed more successful, a finding reported by other researchers[28]. The ordering of directives also impacted how likely the model was to follow them. We found that moving a phrase that was ignored in the coding rationale toward the front of the definition made the model more likely to follow its specifications.
@@
To fully automate the coding process, model output must be reliably readable by a computer. The LLM generates text, which must be interpreted by another script into a data structure, such as a table, for further analysis. Instructing exactly how to format the output produced machine-readable results with GPT-4 and GPT-3.5. Critically, this involves specifying a tag that the interpreting script locates, after which follows a reliably formatted list of codes. Excerpt 1 illustrates how to constrain model output and produce consistent results across queries. Additionally, because GPT tends to be excessively verbose and summarize its output, particularly at higher temperatures, we informed the model that we do not want any output to follow the code list.
@@ 2.2 Selecting a Model and Writing Prompts for Optimal Performance
Studies have shown that LLM decision-making improves when the model is prompted to account for its decisions[29, 30]. This is generally known as chain-of-thought (CoT) prompting or reasoning, and refers to breaking down tasks into specific components, one or more of which involve planning for future steps or reflection on previous ones. Our prompts, which can be viewed in full in the appendix, apply chain-of-thought prompting by including 1) a role assignment step, informing the machine of its purpose, 2) a task description step, specifying the code definition, 3) a justification step, instructing the model to provide a rationale for its decision, and 4) a decision step, wherein the model delivers its ultimate analysis in a consistent, machine-readable format.

### 12 Feb 2024 v2

@@ 1.1 Automating Content Analysis: Past and Present
Early studies of content analysis with LLMs are encouraging. Xiao et al. [9] demonstrate moderate success, Cohen’s $\kappa=0.61$ and $\kappa=0.38$, in two linguistic tasks using GPT-3. Chew et al. [10] report high success on many of 19 tasks across three datasets, and results that are indistinguishable from random for others. It is difficult to evaluate their results due to the choice of Gwet’s AC1, which is biased toward agreement on negative codings rather than positive, whereas most standard measures of intercoder reliability do the opposite [24]. However, we commend Chew et al.’s approach to adapting codebooks for LLMs, which is communicated with great detail and clarity. A survey of 20 empirical pieces reports “mixed-results” of using GPT-3 to automate “text annotation,” a term that ties their framing to “data annotation”, labeling data for in machine learning [25], rather than content analysis in the tradition of grounded theory [26].
We present here three advances to these studies. 1) We report the first methodological account of automating qualitative coding using GPT-4, which, along with other recent models, greatly improves upon many of GPT-3’s capabilities [27, 28, 19]. 2) We provide the first conclusive evidence that LLMs are capable of human-equivalent performance in qualitative coding, and do so on larger passages of text, wherein meaning is often woven through multiple interrelated clauses. 3) We demonstrate that GPT is better at interpreting text when it is tasked with justifying its coding decisions (chain-of-thought prompting) rather than applying codes without an accompanying explanation.
@@ 2.1 Adapting a Codebook for an LLM
Figure 1 demonstrates an effective method of prompting GPT to provide rationale for its code selections. The initial instruction is given by the Justification section of the prompt, and solicited again in the Decision/Formatting box.
Figure 1: The chain-of-thought prompt sequence.
@@
Human coders do not rely solely on a written codebook. Their interpretation of the codes is enriched through the codebook development process, discussions with fellow coders, and supplementary oral instructions. An LLM lacks this interactive and historical context and must interpret codes entirely from written descriptions. Our work modifying the codebook for GPT revealed information that, while implicitly understood by the code developers, wasn’t explicitly stated in the code descriptions. This process not only aided in refining the codebook for automated coding, but also improved our own understandings of the codes. This ultimately led to clearer definition of the codes, thereby enhancing future manual coding processes as well. Figure 2 demonstrates how the Monumental Memorialization and Social/Political Advocacy codes were redefined to improve GPT’s comprehension.
Often, we encountered cases where ambiguous phrasing was obvious to humans, but challenging for the LLM. Our codebook contains two codes that relate to Du Bois’s reputation among academics and activists. These codes are meant to evaluate whether Du Bois appears in a news story because either an academic or activist mentioned him. Initially, we titled this code “Academic Repute,” which worked well for human coders. GPT, however, consistently misinterpretted this code as pertaining to Du Bois’s esteem as an academic, rather than among or by academics (the meaning of “among” remains ambiguous even here). We tried numerous iterations this code without success. Nevertheless, altering the title of the code to the far more literal “Out of the Mouth of Academics” dramatically improved performance, even when paired with the original code description. In another case, the code titled “Social/Political Activism” was revised to “Social/Political Advocacy” (Figure 2, C) because GPT did not consider social critique to be a form of activism, even when it was specifically instructed to.
We found that words indicating how much the model should draw on context or its own outside knowledge had large impacts on the model’s outputs, often to the desired effect. In particular, instructing the model to restrict itself to “explicit” meanings, or to draw on “implicit” meanings, often helped the model with part of a code description it had struggled with. Figure 2 B and D demonstrate the addition of such verbiage to control scope.
Both mandatory (do) and prohibitory (do not) phrasing were observed by the model, though mandatory phrasing seemed more successful, a finding reported by other researchers [29]. The ordering of directives also impacted how likely the model was to follow them. We found that moving a phrase that was ignored in the coding rationale toward the front of the definition made the model more likely to follow its specifications, as in Figure 2 A. When a very specific problem was observed repeatedly, it was sometimes necessary to add a directive to correct it, as in Figure 2 H.
Figure 2: Two examples of prompt redefinition. Colored, alphabetically labeled blocks of text show alterations derived through iterative code refinement. Italics draw attention to direction to constrain interpretive scope to implicit or explicit information.
@@
To fully automate the coding process, model output must be reliably readable by a computer. The LLM generates text, which must be interpreted by another script into a data structure, such as a table, for further analysis. Instructing exactly how to format the output produced machine-readable results with GPT-4 and GPT-3.5. Critically, this involves specifying a tag that the interpreting script locates, after which follows a reliably formatted list of codes. The Decision/Formatting component Figure 1 illustrates how to constrain model output and produce consistent results across queries. Additionally, because GPT tends to be excessively verbose and summarize its output, particularly at higher temperatures, we informed the model that we do not want any output to follow the code list.
@@ 2.2 Selecting a Model and Writing Prompts for Optimal Performance
Studies have shown that LLM decision-making improves when the model is prompted to account for its decisions [30, 31]. This is generally known as chain-of-thought (CoT) prompting or reasoning, and refers to breaking down tasks into specific components, one or more of which involve planning for future steps or reflection on previous ones. Our prompts, which can be viewed in full in the appendix, apply chain-of-thought prompting by including 1) a role assignment step, informing the machine of its purpose, 2) a task description step, specifying the code definition, 3) a justification step, instructing the model to provide a rationale for its decision, and 4) a decision step, wherein the model delivers its ultimate analysis in a consistent, machine-readable format. An example of the chain-of-thought prompt sequence is given by Figure 1.